# Fermionic Averaged Circuit Eigenvalue Sampling

### v1

@@ III Measurement Protocol
Efficiently estimating noise in a quantum system requires that we fit to models with certain simplifying assumptions. Following from definition 8, we assume the parameters describing the noise channel $\mathcal{E}$ are independent of time and, furthermore, depend only on the gate we apply. This is similar to the time-stationary and Markovian assumptions of [22, Def. 2], but unlike Ref. [22], and much related literature, we allow wide latitude for gate dependence. In that regard our protocol can be viewed as a kind of “fermionic gate-set tomography” in the FLO-twirled subspace [6, 23, 7, 8].
@@
We broadly refer to the map that takes ideal physical operations—state preparations, unitary gate operations, and measurements—to their parameterized noisy versions as a model (and we view the FACES model given in definition 1 as a special case of this). Denote the set of idealized physical operations by $\mathcal{S}$ and the set of noisy operations by $\mathcal{N}$, then the model is a bijective map $\phi:\mathcal{S}\rightarrow\mathcal{N}$. We only consider circuits composed of gates for which the model is defined. This is the realizable setting. If the “true” noise belongs to the realizeable setting, then we provide convergence/correctness guarantees that we can learn that model. Otherwise, one of our assumptions is violated. We next present some refining assumptions about our models in the following subsection.
III.1 Model assumptions
@@
Denote a noisy quantum circuit by $\widetilde{\mathcal{U}}=\prod_{g=1}^{L}\widetilde{\mathcal{U}}_{g}$, where each noisy gate factor is given by $\widetilde{\mathcal{U}}_{g}\coloneqq\mathcal{U}_{g}\mathcal{E}_{g}$ as in definition 8 extended to circuits as in eq. 35. The ideal part $\mathcal{U}_{g}$ of each noisy gate is known from the physical operation we applied, but the accompanying noise channel $\mathcal{E}_{g}$ contains free parameters to be learned. In principle, this definition allows for the case where we have multiple ways of physically implementing the same idealized circuit, resulting in multiple circuits with the same ideal part but possibly distinct accompanying noise channels. As stated above, we will restrict to the setting where the applied noise channel is unique to the corresponding idealized gate (that is, our model $\phi$ is bijective), but it is straightforward to relax this assumption. In addition, the noisy part of each gate may also depend on the idealized gate itself. Under these assumptions, the FACES model given in definition 1 is a special case of the general model described in this section where each individual FLO unitary gate is associated to its FLO-twirled noisy version through its collection of gate eigenvalues. In summary, we assume that any noise model we wish to learn is Markovian in at least two senses: first, that it can be described by a completely positive trace-preserving map, but also that the noise specific to a given gate depends only on the idealized gate itself.
@@
$||\mathcal{I}-\mathcal{E}_{j}^{\mathds{F}(2n)}||_{\infty}\leq c^{\prime}$ (39)
for all noise channels $\mathcal{E}_{j}$ that we are trying to learn. Here again, $c^{\prime}$ is a sufficiently small constant. Since our channels are always diagonal in the Pauli basis, this is equivalent to requiring that any noisy-channel eigenvalue lies in the interval $[1-c^{\prime},1]$ (in the proof of theorem 13, we require that $c^{\prime}=\frac{1}{2}$). In practice, we routinely expect that these numbers $c,c^{\prime}$ are less than $0.05$ or even $0.01$ in some quantum architectures, so these are not strong assumptions.
III.2 FLO-Specific Considerations
A fundamental challenge attendant to the setting of noisy FLO circuits arises due to the fact that the group $\mathds{F}(2n)$ is continuous. Since we can only ever estimate finitely many noise parameters, we thus assume that the noise can nonetheless be described by a finite set of channels. For completeness, we choose a particular gate set given by $\mathcal{G}\coloneqq\{e^{i\theta Z_{j}}\}_{j,\theta}\cup\{G_{j}(H,H)\}_{j}$ in the qubit picture, where $\theta\in[0,2\pi)$ is allowed to be an arbitrary continuous parameter for each $j$. The matrix $H$ is the single-qubit Hadamard operator, and $G_{j}(A,B)$ is the 2-qubit matchgate as defined in section II.1. While the 2-qubit gates $\{G_{j}(H,H)\}_{j}$ form a discrete set, the single qubit rotations $\{e^{i\theta Z_{j}}\}_{j,\theta}$ do not. We partition the interval $[0,2\pi)$ into $N$ bins of the form $B_{k}\coloneqq[\frac{2\pi(k-1)}{N},\frac{2\pi k}{N})$ for $k\in[N]$, and we assume that the single-qubit rotation $e^{i\theta Z_{j}}$ is subject to a fixed noise channel $\mathcal{E}_{1,jk}$ if $\theta\in B_{k}$. Similarly, we assume that each of the $\{G_{j}(H,H)\}_{j}$ is subject to a fixed noise channel $\mathcal{E}_{2,j}$ An interesting related question is whether we can use these methods to learn a finite set of noise generators for a continuous family of noise operations. We leave this question for future work, and we return to it in the Discussion.
@@ III.2 FLO-Specific Considerations
III.3 Protocol Details
Our measurement protocol for obtaining FLO-averaged circuit eigenvalues closely follows the ACES protocol of Ref. [9], and it is detailed graphically in fig. 1. To learn all the eigenvalues of a FLO-twirled noise channel, we present two complementary protocols, which differ slightly in the required input state, measurement basis, and net FLO circuit applied. An important unitary conjugation channel in our analysis is $\mathcal{U}_{+}(X)=U_{+}XU_{+}^{\dagger}$, defined in [17] to be a channel satisfying
@@ III.3 Protocol Details
where $r\in\{``x",``z"\}$, and the net ideal unitary varies with $r$ as follows: for all $\widetilde{\mathcal{U}}^{(j)}\in\mathcal{C}_{z}$, we have $\mathcal{U}^{(j)}=\mathcal{I}$ (the identity channel), and for all $\widetilde{\mathcal{U}}^{(j)}\in\mathcal{C}_{x}$ we have $\mathcal{U}^{(j)}=\mathcal{U}_{+}$. Recall that a tilde is used to indicate the noisy verion of a gate or circuit.
We can organize our circuits into two cases depending on whether the first qubit is prepared in the $x$ or $z$ basis; we call these “type-$x$” and “type-$z$”, respectively. The readout data, and thus the expected signal, takes two different forms in these two cases. We thus organize our empirical estimates of the outcome probabilities according to the circuit (or equivalently, readout) type. Readout data of type-$z$ is the Hamming weight of an $n$-bit string, whereas readout data of type-$x$ consists of a sign ($\pm$) together with the Hamming weight of an $(n-1)$-bit string. When we collect these data into empirical probability estimates over the relevant Hamming weights, we denote these empirical estimates by $\mathbf{P}$ regardless of the $x$/$z$ type. Nevertheless, we can distinguish the type by adding an index $s$ that takes values in $\{+,-,0\}$ where 00 indicates $z$-type data, and $\pm$ corresponds to the sign label of $x$-type data. For example, the estimated probability of weight-$j$ in a $+$-outcome $x$-type experiment would be labeled $\mathrm{P}_{+,j}$, and so on.
@@
for $\widetilde{\mathcal{U}}^{(j)}\in\mathcal{C}_{z}$. Similarly, define the $x$-type Born-rule probability distribution in experiment $j$ as
@@
The Born-rule probability distributions are related to the circuit eigenvalues of a FLO-twirled quantum circuit. Let $\Lambda^{(j)}_{k}$ denote the degree-$k$ eigenvalue of circuit $j$ in either ensemble $\mathcal{C}_{x}$ or $\mathcal{C}_{z}$, for $k\in[2n]$. We collect these eigenvalues into vectors as
@@
For an ensemble of noisy FLO circuits $\mathcal{C}_{r}$,
@@
$\begin{cases}\mathbf{V}_{z}=\bigoplus_{j\in\mathcal{C}_{z}}\left[\mathbf{M}^{(n)}\cdot(\mathbf{d}^{(n)})^{-1}\right]&\\ \mathbf{V}_{x}=\bigoplus_{j\in\mathcal{C}_{x}}\left\{\left[\mathbf{M}^{(n-1)}\cdot(\mathbf{d}^{(n-1)})^{-1}\right]\oplus\left[\mathbf{M}^{(n-1)}\cdot(\mathbf{d}^{(n-1)})^{-1}\right]\right\}\cdot(\sqrt{2}\mathbf{H}\otimes\mathbf{I}_{n})&\end{cases}$ (53)
@@
where the design matrix element $A_{j,g}$ corresponds to the number of occurences of gate $g$ in circuit $j$. Physically, we sample the born-rule probabilities defined in eq. 43 and eq. 44 to obtain empirical estimates $\hat{\mathbf{P}}^{(j)}$ of these distributions for each circuit $j$, and we apply the transformations in lemma 11 to produce estimators $\hat{\mathbf{\Lambda}}^{(j)}$ for the associated circuit eigenvalues. Finally, we fit our gate eigenvalue estimates to our circuit eigenvalue data using a log-linear model. That is, our estimator for each Majorana degree $k$ is the solution to a system of equations of the form $\mathbf{Ax}=\mathbf{b}$, where $\mathbf{A}$ is the design matrix, $x_{g,k}=-\mathrm{log}(\xi_{g,k})$, and $b_{j,k}=-\mathrm{log}(\Lambda^{(j)}_{k})$. More formally, we give the following definition
@@
That is, we take the estimator given by ordinary least-squares, projected onto the positive orthant for each Majorana degree $k$. One could of course consider other estimators, such as $\hat{\mathbf{x}}^{\prime}$ taken to be the element of the positive orthant that minimizes $\lVert\mathbf{A}\hat{\mathbf{x}}^{\prime}-\hat{\mathbf{b}}\rVert_{2}$, or even an estimator that maximizes a likelihood function subject to a positivity constraint. These are convex programs and would presumably be efficient to consider as practical estimators. However, our choice of estimator is convenient from the standpoint of proving sample-complexity guarantees. We give a pseudocode description of the relevant subroutines in appendix A.
III.4 Sample Complexity
In this section, we analyze the sample complexity required to bound the error on the estimator $\hat{\boldsymbol{\xi}}$ in eq. 55. We provide the following theorem:
@@ III.4 Sample Complexity
We thus require $S=O\left(\frac{n\lVert\mathbf{A}^{+}\rVert_{\infty}^{2}\log(m/\delta)}{\epsilon^{2}}\right)$ samples per experiment to guarantee error $\epsilon$ with probability $1-\delta$.
@@
(lemma 14) The logarithms of the circuit and gate eigenvalues satisfy a linear relationship
@@
for each $i$, such that $u_{i}(0)=\xi_{i}$ and $u_{i}(1)=\hat{\xi}_{i}$. Applying lemma 15 to each such function separately,
@@
From the second to the third line above, we replaced the maximum over indices $i$ in $\lVert\cdot\rVert_{\infty}$ with a maximization first over each experiment and then within each experiment. We then applied the norm inequality $\lVert\cdot\rVert_{\infty}\leq\lVert\cdot\rVert_{1\rightarrow\infty}\lVert\cdot\rVert_{1}$ experiment by experiment and used the fact that the maximum absolute element of $\mathbf{V}$ is 1 in each case from eq. 53. Combining eq. 80 with eq. 70 gives the result. ∎
IV Numerics
Figure 2: Histogram of relative errors for $102n=$ 10 fermionic modes with $1000$ circuits for each experiment type. See main text for description.
Finally, we demonstrate the efficacy of our protocol numerically in fig. 2, where we plot a histogram of relative gate-eigenvalue errors for $n=5$ qubits ($102n=$ 10 fermionic modes) and $m=1000$ circuits for each experiment type. Our noise model is given by a random $2$-qubit Pauli channel following each gate we apply, where the Pauli channel probabilities are chosen uniformly from the interval $[10^{-2}-10^{-3},10^{-2}+10^{-3}]$. Columns are partitioned according to $x$- and $z$-type estimates, and rows are partitioned according to whether the different histograms correspond to different Majorana weight-$k$ or different numbers $S$ of samples from the Born-rule probability distribution (we assume that we can apply FLO-twirled circuits exactly). We also partition $Z$-rotation angles into 46 bins of width $\Delta\theta=\frac{\pi}{23}$. We observe that the protocol is able to routinely achieve less than 5% relative error for most of the eigenvalues estimated, and that this steadily improves with the number of shots.
V Conclusion
We have shown how to learn the matchgate-averaged noise parameters on many matchgates simultaneously via FACES, our fermionic adaptation of the ACES protocol. We envision that this will enable experiments to build a complete characterization of averaged noise parameters for universal gate sets, though several open problems remain. Foremost is the technical challenge of performing matchgate twirling. Unlike Pauli twirling, it seems unlikely that matchgate twirling can be achieved in constant depth. However, there have recently been promising results to show that low-depth circuits can nevertheless approximate Haar random unitaries [24, 25]. If a corresponding result holds for Haar random matchgates, then this issue can be mitigated, if not entirely avoided. Alternatively, we might consider only reproducing the Haar ensemble of FLO circuits up to the second moment, which is sufficient for exactly twirling the channel. It is known that this can be achieved using only gates in the intersection of the matchgate and Clifford groups [26, 27].

### v2

@@ 3 Measurement Protocol
Efficiently estimating noise in a quantum system requires that we fit to models with certain simplifying assumptions. Following from definition˜8, we assume the parameters describing the noise channel $\mathcal{E}$ are independent of time and, furthermore, depend only on the gate we apply. This is similar to the time-stationary and Markovian assumptions of [23, def. 2], but unlike ref. [23], and much related literature, we allow wide latitude for gate dependence. In that regard our protocol can be viewed as a kind of “fermionic gate-set tomography” in the FLO-twirled subspace [6, 24, 7, 8].
@@
We broadly refer to the map that takes ideal physical operations—state preparations, unitary gate operations, and measurements—to their parameterized noisy versions as a model (and we view the FACES model given in definition˜1 as a special case of this). Denote the set of idealized physical operations by $\mathcal{S}$ and the set of noisy operations by $\mathcal{N}$, then the model is a bijective map $\phi:\mathcal{S}\rightarrow\mathcal{N}$. We only consider circuits composed of gates for which the model is defined. This is the realizable setting. If the “true” noise belongs to the realizable setting, then we provide convergence/correctness guarantees that we can learn that model. Otherwise, one of our assumptions is violated. We next present some refining assumptions about our models in the following subsection.
3.1 Model assumptions
@@
Denote a noisy quantum circuit by $\widetilde{\mathcal{U}}=\prod_{g=1}^{L}\widetilde{\mathcal{U}}_{g}$, where each noisy gate factor is given by $\widetilde{\mathcal{U}}_{g}\coloneqq\mathcal{U}_{g}\mathcal{E}_{g}$ as in definition˜8, extended to a circuit of length $L$ as in eq.˜35. The ideal part $\mathcal{U}_{g}$ of each noisy gate is known from the physical operation we applied, but the accompanying noise channel $\mathcal{E}_{g}$ contains free parameters to be learned. In principle, this definition allows for the case where we have multiple ways of physically implementing the same idealized circuit, resulting in multiple circuits with the same ideal part but possibly distinct accompanying noise channels. As stated above, we will restrict to the setting where the applied noise channel is unique to the corresponding idealized gate (that is, our model $\phi$ is bijective), but it is straightforward to relax this assumption. In addition, the noisy part of each gate may also depend on the idealized gate itself. In principle, the twirling gates $\mathcal{U}^{\dagger}$ and $\mathcal{V}\mathcal{U}\mathcal{V}^{\dagger}$ in eq.˜27 are noisy in the FLO-twirling case as well, but we assume that this noise is only weakly correlated with $\mathcal{V}$, since we can implement them without necessarily implementing $\mathcal{V}$ due to the FLO group structure. Our expectation is that, in practice, this noise will introduce some additional uncertainty on our gate-eigenvalue estimates. Under these assumptions, the FACES model given in definition˜1 is a special case of the general model described in this section where each individual FLO unitary gate is associated to its FLO-twirled noisy version through its collection of gate eigenvalues. In summary, we assume that any noise model we wish to learn is Markovian in at least two senses: first, that it can be described by a completely positive trace-preserving map, but also that the noise specific to a given gate depends only on the idealized gate itself.
@@
[…]
for all noise channels $\mathcal{E}_{j}$ that we are trying to learn. Here again, $c^{\prime}$ is a sufficiently small constant. Since our channels are always diagonal in the Pauli basis, this is equivalent to requiring that any noisy-channel eigenvalue lies in the interval $1-c^{\prime},1]$ (in the proof of theorem˜13, we require that $c^{\prime}=\frac{1}{2}$). In practice, we routinely expect that these numbers $c,c^{\prime}$ are less than $0.05$ or even $0.01$ in some quantum architectures, so these are not strong assumptions.
3.2 FLO-Specific Considerations
A fundamental challenge attendant to the setting of noisy FLO circuits arises due to the fact that the group $\mathds{F}(2n)$ is continuous. Since we can only ever estimate finitely many noise parameters, we thus assume that the noise can nonetheless be described by a finite set of channels. For completeness, we choose a particular gate set given by $\mathcal{G}\coloneqq\{e^{i\theta Z_{j}}\}_{j,\theta}\cup\{G_{j}(H,H)\}_{j}$ in the qubit picture, where $\theta\in[0,2\pi)$ is allowed to be an arbitrary continuous parameter for each $j$. The matrix $H$ is the single-qubit Hadamard operator, and $G_{j}(A,B)$ is the 2-qubit matchgate as defined in section˜2.1. While the 2-qubit gates $\{G_{j}(H,H)\}_{j}$ form a discrete set, the single qubit rotations $\{e^{i\theta Z_{j}}\}_{j,\theta}$ do not. We partition the interval $0,2\pi)$ into $N$ bins of the form $B_{k}\coloneqq[\frac{2\pi(k-1)}{N},\frac{2\pi k}{N})$ for $k\in[N]$, and we assume that the single-qubit rotation $e^{i\theta Z_{j}}$ is subject to a fixed noise channel $\mathcal{E}_{1,jk}$ if $\theta\in B_{k}$. Similarly, we assume that each of the $\{G_{j}(H,H)\}_{j}$ is subject to a fixed noise channel $\mathcal{E}_{2,j}$ An interesting related question is whether we can use these methods to learn a finite set of noise generators for a continuous family of noise operations. We leave this question for future work, and we return to it in the Discussion.
@@ 3.2 FLO-Specific Considerations
3.3 Protocol Details
Our measurement protocol for obtaining FLO-averaged circuit eigenvalues closely follows the ACES protocol of ref. [9], and it is detailed graphically in fig.˜1. To learn all the eigenvalues of a FLO-twirled noise channel, we present two complementary protocols, which differ slightly in the required input state, measurement basis, and net FLO circuit applied. An important unitary conjugation channel in our analysis is $\mathcal{U}_{+}(X)=U_{+}XU_{+}^{\dagger}$, related to the unitary $U(R_{+})$ defined in [18], and satisfying
@@ 3.3 Protocol Details
where $r\in\{``x",``z$"\}, and the net ideal unitary varies with $r$ as follows: for all $\widetilde{\mathcal{U}}^{(j)}\in\mathcal{C}_{z}$, we have $\mathcal{U}^{(j)}=\mathcal{I}$ (the identity channel), and for all $\widetilde{\mathcal{U}}^{(j)}\in\mathcal{C}_{x}$ we have $\mathcal{U}^{(j)}=\mathcal{U}_{+}$. Recall that a tilde is used to indicate the noisy version of a gate or circuit. To avoid overloading the superscript notation, we denote $\mathcal{C}_{r}^{\mathds{F}(2n)}$ as the ensemble whose elements correspond to the FLO-twirled elements of $\mathcal{C}_{r}$, and we drop the explicit superscript on the elements of $\mathcal{C}_{r}^{\mathds{F}(2n)}$ themselves, with the understanding that they are twirled.
We can organize our circuits into two cases depending on whether the first qubit is prepared in the $x$ or $z$ basis; we call these “type-xx” and “type-zz”, respectively. The readout data, and thus the expected signal, takes two different forms in these two cases. We thus organize our empirical estimates of the outcome probabilities according to the circuit (or equivalently, readout) type. Readout data of type-$z$ is the Hamming weight of an $n$-bit string, whereas readout data of type-$x$ consists of a sign ($\pm$) together with the Hamming weight of an (n1)(n-1)-bit string. When we collect these data into empirical probability estimates over the relevant Hamming weights, we denote these empirical estimates by $\mathbf{P}$ regardless of the xx/$z$ type. Nevertheless, we can distinguish the type by adding an index $s$ that takes values in $\{+,-,0\}$ where 0 indicates $z$-type data, and $\pm$ corresponds to the sign label of $x$-type data. For example, the estimated probability of weight-$j$ in a ++-outcome $x$-type experiment would be labeled $\mathrm{P}_{+,j}$, and so on.
@@
for $\widetilde{\mathcal{U}}^{(j)}\in\mathcal{C}_{z}^{\mathds{F}(2n)}$. Similarly, define the $x$-type Born-rule probability distribution in experiment $j$ as
@@
The Born-rule probability distributions are related to the circuit eigenvalues of a FLO-twirled quantum circuit. Let $\Lambda^{(j)}_{k}$ denote the degree-$k$ eigenvalue of circuit $j$ in either ensemble $\mathcal{C}_{x}^{\mathds{F}(2n)}$ or $\mathcal{C}_{z}^{\mathds{F}(2n)}$, for $k\in[2n]$. We collect these eigenvalues into vectors as
@@
For an ensemble of noisy and FLO-twirled FLO circuits $\mathcal{C}^{\mathds{F}(2n)}_{r}$,
@@
[…]
@@
where the design matrix element $A_{j,g}$ corresponds to the number of occurences of gate $g$ in circuit $j$. Physically, we sample the Born rule probabilities defined in eq.˜43 and eq.˜44 to obtain empirical estimates $\hat{\mathbf{P}}^{(j)}$ of these distributions for each circuit $j$, and we apply the transformations in lemma˜11 to produce estimators $\hat{\mathbf{\Lambda}}^{(j)}$ for the associated circuit eigenvalues. Finally, we fit our gate eigenvalue estimates to our circuit eigenvalue data using a log-linear model. That is, our estimator for each Majorana degree $k$ is the solution to a system of equations of the form $\mathbf{Ax}=\mathbf{b}$, where $\mathbf{A}$ is the design matrix, $x_{g,k}=-\mathrm{log}(\xi_{g,k})$, and $b_{j,k}=-\mathrm{log}(\Lambda^{(j)}_{k})$. More formally, we give the following definition
@@
That is, we take the estimator given by ordinary least-squares, projected onto the positive orthant for each Majorana degree $k$. One could of course consider other estimators, such as $\hat{\mathbf{x}}^{\prime}$ taken to be the element of the positive orthant that minimizes $\lVert\mathbf{A}\hat{\mathbf{x}}^{\prime}-\hat{\mathbf{b}}\rVert_{2}$, or even an estimator that maximizes a likelihood function subject to a positivity constraint. These are convex programs and would presumably be efficient to consider as practical estimators. However, our choice of estimator is convenient from the standpoint of proving sample-complexity guarantees. We give a pseudocode description of the relevant subroutines in appendix˜A.
3.4 Sample Complexity
In this section, we analyze the sample complexity required to bound the error on the estimator $\hat{\boldsymbol{\xi}}$ in eq.˜55. We provide the following theorem:
@@ 3.4 Sample Complexity
We thus require $S=O\left(\frac{n\lVert\mathbf{A}^{+}\rVert_{\infty}^{2}\log(J/\delta)}{\epsilon^{2}}\right)$ samples per experiment to guarantee error $\epsilon$ with probability $1-\delta$.
@@
(lemma˜14) The logarithms of the circuit and gate eigenvalues satisfy a linear relationship
@@
for each $i$, such that $u_{i}(0)=\xi_{i}$ and $u_{i}(1)=\hat{\xi}_{i}$. Applying lemma˜15 to each such function separately,
@@
From the second to the third line above, we replaced the maximum over indices $i$ in $\lVert\cdot\rVert_{\infty}$ with a maximization first over each experiment and then within each experiment. We then applied the norm inequality $\lVert\cdot\rVert_{\infty}\leq\lVert\cdot\rVert_{1\rightarrow\infty}\lVert\cdot\rVert_{1}$ experiment by experiment and used the fact that the maximum absolute element of $\mathbf{V}$ is 1 in each case from eq.˜53. Combining eq.˜80 with eq.˜70 gives the result. ∎
4 Numerics
Figure 2: Histogram of relative errors for $2n=10$ fermionic modes with $1000$ circuits for each experiment type. See main text for description.
Finally, we demonstrate the efficacy of our protocol numerically in fig.˜2, where we plot a histogram of relative gate-eigenvalue errors for $n=5$ qubits ($2n=10$ fermionic modes) and $m=1000$ circuits for each experiment type. Our noise model is given by a random $2$-qubit Pauli channel following each gate we apply, where the Pauli channel probabilities are chosen uniformly from the interval $10^{-2}-10^{-3},10^{-2}+10^{-3}]$. Columns are partitioned according to $x$- and $z$-type estimates, and rows are partitioned according to whether the different histograms correspond to different Majorana weight-$k$ or different numbers $S$ of samples from the Born-rule probability distribution (we assume that we can apply FLO-twirled circuits exactly). We also partition $Z$-rotation angles into 46 bins of width $\Delta\theta=\frac{\pi}{23}$. We observe that the protocol is able to routinely achieve less than 5% relative error for most of the eigenvalues estimated, and that this steadily improves with the number of shots.
5 Conclusion
We have shown how to learn the matchgate-averaged noise parameters on many matchgates simultaneously via FACES, our fermionic adaptation of the ACES protocol. We envision that this will enable experiments to build a complete characterization of averaged noise parameters for universal gate sets, though several open problems remain. Foremost is the technical challenge of performing matchgate twirling. Unlike Pauli twirling, it seems unlikely that matchgate twirling can be achieved in constant depth. However, there have recently been promising results to show that low-depth circuits can nevertheless approximate Haar random unitaries [25, 26]. If a corresponding result holds for Haar random matchgates, then this issue can be mitigated, if not entirely avoided. Alternatively, we might consider only reproducing the Haar ensemble of FLO circuits up to the second moment, which is sufficient for exactly twirling the channel. It is known that this can be achieved using only gates in the intersection of the matchgate and Clifford groups [27, 28]. Finally, we might employ restricted circuit structure to address this challenge in a similar strategy to interleaved randomized benchmarking (IRB) [29], which has historically enjoyed a great deal of success in practice. In a setting where the circuit structure is limited to the FLO analog of IRB, for example, we expect that FACES will exhibit a “repeated convolution structure” similar to that of matchgate benchmarking [18], since we are considering Haar averages over the group of FLO unitaries.